AutoWrecks, Inc. sells auto insurance. AutoWrecks keeps close tabs on its customers driving records, updating its rates according to the trends indicated by these records. AutoWrecks records indicate that, in a typical year, roughly 70% of the companys customers do not commit a moving violation, 10% commit exactly one moving violation, 15% commit exactly two moving violations, and 5% commit three or more moving violations. This past years driving records for a random sample of 100 Auto Wrecks customers is summarized in the first row of numbers in Table 1 below. This row gives this years observed frequencies for each moving violation category for the sample of 100 AutoWrecks customers. The second row of numbers gives the frequencies expected for a sample of 100 AutoWrecks customers if the moving violations distribution for this year is the same as the distribution for a typical year. The bottom row of numbers in Table 1 contains the values BO (fo-fr)2 (Jo J (Observed frequency - Expected frequency)2 Expected frequency for each of the moving violation categories. Fill in the missing values of Table 1. Then, using the 0.10 level of significance, perform a test of the hypothesis that there is no difference between this years moving violation distribution and the distribution in a typical year. Then complete Table 2. (fo-fe2 Round your responses for the expected frequencies in Table 1 to at least two decimal places. Round your in table to ease no decina puces. Round your opin ve - responses in Table 1 to at least three E decimal places. Round your responses in Table 2 as specified. Ei send data ed to Excel Table 1: Information about the SamplePlease answer my problem ASAP

Thank you.
Answer #1

THe missing values are Table 1 are as follows

No Violations Exactly one violation Exactly two violations Three or more Violations Total
Observed Frequency (fO) 58 13 21 8 100
Expected Frequency (fE) 70 10 15 5 100
(fO - fE)2/fE 2.057 0.900 2.400 1.800 7.157

The type of the test statistic is Chi-square ( \chi ^{2}) test for goodness of fit

The value of the test statistic is 7.157 \chi ^{2}=\sum \frac{(f_{O}-f_{E})^{2}}{f_{E}}=7.157

The critical value of the Chi-Square  ( \chi ^{2}) test at 0.10 level of significance and (4-1)=3 df is   \chi ^{2}=6.25

Since the test statisti value is greater than the critical value, i.e., 7.157>6.25, we can reject the hypothesis that there is no difference between this year's moving violation distribution and the distribution in a "typical" year at 0.10 significance lvel

The correct answer is YES